carbonyl bromide, COBr2, decomposes to CO and Br2 with an Kc= .190 . A .015 mole sample of carbonyl bromide is place in a 2.5 L flask. what will be the equilibrium concentrations of the reaction?
Answer #1

Consider a equilibrium reaction, COBr 2 CO + Br 2

For above reaction, K c = [ CO ] [ Br 2 ] / [ COBr 2 ] = 0.190

Let's use ICE table.

M COBr 2 CO + Br 2
I 0.006
C -X +X +X
E 0.006 - X X X

( [ COBr 2 ] = 0.015 mol / 2.5 L = 0.006 M )

K c = ( X) (X) / 0.006 - X = 0.190

X 2 / 0.006 - X = 0.190

X 2 = 0.006 - X ( 0.190 )

X 2 = 0.00114 - 0.190 X

X 2 + 0.190 X - 0.00114= 0

Comparing above equation with a X 2 + b X +c = 0 , we get a = 1 , b = 0.190 and c = - 0.00114

We know that, X = - b +/- b 2 -4 ac / 2a

X = - 0.190 +/- ( 0.190 ) 2 - 4 (1) ( - 0.00114 ) / 2 (1)

X = - 0.1958 or X = 0.005822

We can choose X = 0.005822

Equilibrium concentration of CO = X = 0.005822 M

Equilibrium concentration of Br 2 = X = 0.005822 M

Equilibrium concentration of COBr 2 = 0.006 - X = 0.000178 M